Boltzmann’s constant is 1.38066×10−23 J/K,
and the universal gas constant is
8.31451 J/K · mol.
If 1.6 mol of a gas is confined to a 5.2 L
vessel at a pressure of 3.9 atm, what is the average kinetic energy of a gas molecule?
Answer in units of J.
1
Expert's answer
2020-03-13T10:53:54-0400
"E=\frac{3}{2}kT"

"pV=nRT"

"E=\frac{3pVk}{2nR}"

"E=\frac{3(1.38066\cdot 10^{-23})(3.9\cdot 101325)(5.2\cdot 10^{-3})}{2(8.31451)(1.6)}"

"E=3.2\cdot 10^{-21}\ J"